Opened 7 years ago

Closed 14 months ago

#469 closed flight processing (fixed)

ET12/17, flight day 320/2012, Corbetti

Reported by: mark1 Owned by: benj
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anhi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/ethiopia/ET12_17-2012_320_Corbetti/

Data arrived from ARSF via SATA disk on 06/12/2012

ET12/17 Scientific objective:

ET12/17 Priority: a4h

PI: David Pyle

Sensors:

Camera (05/03/2013)
Eagle (Requested)
Hawk (Requested)
LiDAR (05/03/2013)

Change History (32)

comment:1 Changed 7 years ago by mark1

GPS basestation data acquired via ftp 17/12/12

comment:2 Changed 7 years ago by tipo

Started Hyperspectral Processing.
used ppp for navigation.

comment:3 Changed 7 years ago by tipo

Finished Hyperspectral Processing

comment:4 Changed 7 years ago by anhi

Hyperspectral Delivery check:

  • Two xml files in the project information folder, named differently but look the same.
  • Line e11 join with e12 don't line up.
  • Mention in Read Me the holes in the lines caused by cloud.

comment:5 Changed 7 years ago by edfi

Navigation Re-processed using the new files and LiDAR started

comment:6 Changed 7 years ago by anhi

RCD Delivery Created.

comment:7 Changed 7 years ago by edfi

LiDAR delivery created awaiting D/C

comment:8 Changed 7 years ago by tipo

Checked over lines e11 and e12 again, They don't line up with each other exactly, but there doesn't seem to be an sct value that matches up the join any better without causing distortion.

comment:9 Changed 7 years ago by tipo

Checked camera delivery, no problems found.

comment:10 Changed 7 years ago by knpa

Beginning LiDAR DC

comment:11 Changed 7 years ago by anhi

Second Hyperspectral Delivery check

No problems found in changes made.

comment:12 Changed 7 years ago by edfi

Checking single point PPP for Nav

comment:13 Changed 7 years ago by edfi

Single point PPP needed to be adjusted. Reprocessed Nav and re-ran delivery. Will update ReadMe once complete

comment:14 Changed 7 years ago by edfi

ReadMe adjusted. Ready for re-check

comment:15 Changed 7 years ago by knpa

LiDAR delivery check complete

Looks good apart from the following:

  1. In ReadMe, elevation difference between lines should have the line names from logsheet. Also appears to be one missing.
  1. Minor thing but should be gap between comma and Z in "Min X,Y,Z:" in flightline statistics in ReadMe. Can fix with a find and replace in the tex file (someone should also check the script that inserts this).
  1. Should probably mention that line 12 in particular has almost no data (due to cloud?)
  1. Too many false positives on the classification on the lines I checked (4 & 9). Try running las_classify with less strict -isolated setting or omit entirely

comment:16 Changed 7 years ago by edfi

starting to make adjustments

comment:17 Changed 7 years ago by edfi

Adjustments made:

Renamed flightlines to reflect logsheet.

Had a look at X,Y,Z in latex script but it is consistent between letters (i.e. no odd spaces). Tried adding a space before z but made the reverse problem.

Added comment about line 12

Loosened use for -isolated

Created new delivery to update DEM, classification statistics etc

Ready for re-check

comment:18 Changed 7 years ago by knpa

Re-checked and looks good, ready to delivery.

comment:19 Changed 7 years ago by benj

PI has requested that this flight be processed using the December 2012 calibration - please process the hyperspectral data accordingly.

comment:20 Changed 7 years ago by anhi

Delivered LiDAR and rcd data on 05/03/2013 on disk 152 to David Pyle, University of Oxford.

comment:21 Changed 7 years ago by anhi

  • Description modified (diff)

comment:22 Changed 7 years ago by tipo

Reprocessed hyperspectral with new calibration.

comment:23 Changed 7 years ago by edfi

Starting D/C of hyperspectral

comment:24 Changed 7 years ago by edfi

Missing screenshots for a couple of lines, these have been created and included.

Need to add project_information and update ReadMe to include new Screenshots

No other problems

comment:25 Changed 7 years ago by tipo

Created xml file, problem was caused by the FODIS hdr files somehow being in the level1b directory

comment:26 Changed 7 years ago by edfi

Rechecked, all looks ok, ready to deliver

comment:27 Changed 7 years ago by tipo

Renamed posatt/ipas_honeywell/extract/2012230_RTMsg.txt to 2012_230_RTMsg.txt as requested by proj_tidy.sh

comment:28 Changed 7 years ago by tipo

Delivered hyperspectral data to David Pyle, University of Oxford 3/5/2013

comment:29 Changed 7 years ago by tipo

archiving

comment:30 Changed 7 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Owner set to benj

comment:31 Changed 6 years ago by tipo

Made avalable for archiving 13/01/2014

comment:32 Changed 14 months ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.